Can you provide some assist in helping solve this problem...instructions...solve the system by addition or substitution.
4x - 4y = 24
y = -6 + x

Answer by stanbon(48568)

solve the system by addition or substitution.
4x - 4y = 24
y = -6 + x
----------------
Substitute y=-6+x into the 1st equation, as follows:
4x - 4(-6+x) = 24
4x +24 -4x = 24
4x-4x = 24-24
0=0
-----------
This means the two equations are really the same.
The solution for the system of equations is all the points that satisfy
y=-6+x; this is a line with an infinite number of points.
Cheers,
Stan H.